ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 182, 2011

Arsovski, B. A proof of Snevily’s conjecture  505

Beanland, K. An ordinal indexing on the space of strictly singular operators

47

Betancor, J. J., J. C. Fariña, L. Rodríguez-Mesa, A. Sanabria and J. L. Torrea Lusin type and cotype for Laguerre gfunctions

1

Biró, A. A relation between triple products of weight 0 and weight \frac{1}{2} cusp forms

61

Breuil, C. Diagrammes de Diamond et (φ, Γ)-modules  349

Bukh, B., J. Matoušek and G. Nivasch Lower bounds for weak epsilon-nets and stair-convexity

199

Bunimovich, L. A. and A. Yurchenko Where to place a hole to achieve a maximal escape rate

229

Di Scala, A. J. and T. Leistner Connected subgroups of SO(2, n) acting irreducibly on ℝ^{2,n}

103

Fariña, J. C. See J. J. Betancor Feng, T., Z. W. Sun and Q. Xiang Exterior algebras and two

conjectures on finite abelian groups

425

Garbit, R. A note on Furstenberg’s filtering problem  333

Hovey, M. and K. Lockridge The ghost and weak dimensions of rings and ring spectra

31

Izhakian, Z. and L. Rowen Supertropical matrix algebra  383
Kalaj, D. Harmonic maps between annuli on Riemann surfaces  123

Kozlov, D. N. Moduli spaces of tropical curves of higher genus with marked points and homotopy colimits

253

Leistner, T. See A. J. Di Scala Lockridge, K. See M. Hovey Matoušek, J. See B. Bukh

Merimovich, C. Extender-based Magidor–Radin forcing  439

Mishchenko, S. P., V. M. Petrogradsky and A. Regev Characterization of non-matrix varieties of associative algebras

337

Neal, M. and B. Russo Existence of contractive projections on preduals of JBW^{∗}-triples

293

Nivasch, G. See B. Bukh

Petersen, L. and H. Süss Torus invariant divisors  481

Petrogradsky, V. M. See S. P. Mishchenko Regev, A. See S. P. Mishchenko Rodríguez-Mesa, L. See J. J. Betancor Rowen, L. See Z. Izhakian Russo, B. See M. Neal Süss, H. See L. Petersen Sanabria, A. See J. J. Betancor Shelah, S. and A. Usvyatsov Model theoretic stability and cate-

goricity for complete metric spaces

157

Shumyatsky, P. Commutators in residually finite groups  149

Sun, Z. W. See T. Feng Torrea, J. L. See J. J. Betancor Usvyatsov, A. See S. Shelah Xiang, Q. See T. Feng Yurchenko, A. See L. A. Bunimovich